OpenStudy (anonymous):

The length of the hypotenuse of a 30-60-90 triangle is 9. What is the perimeter?

OpenStudy (anonymous):



OpenStudy (anonymous):

the answer is supposed to be in equation form. how would I do that?

OpenStudy (anonymous):

first find x

OpenStudy (anonymous):

okay

OpenStudy (wolf1728):

a 30 60 90 triangle has sides in these proportions 1, sq root (3) and 2 So the sides would be 4.5, sq root (3) * 4.5 and 9 = 4.5, 7.794 and 9 perimeter = 21.294

OpenStudy (anonymous):

x=7.794

OpenStudy (anonymous):

how would I put it in equation form?

OpenStudy (wolf1728):

In equation form sides of a 30 60 90 triangle are in the proportion hyp/2 ; (hyp/2) * sq root of 3 ; hypotenuse *1 9/ 2 ; ((9/2) * 1.732) ; 9

OpenStudy (anonymous):

the answer choices are: A. 27/2+9/2√3 B.9/2+27/2√3 C. 9+27√3 D. 27+9√3

OpenStudy (anonymous):

which one of them would it be?

OpenStudy (anonymous):

i think it might be B..?

OpenStudy (wolf1728):

Okay Perimeter = 9/ 2 + ((9/2) * sq root(3) ) + 9 Still working on those choices

OpenStudy (anonymous):

I still think it's B...

OpenStudy (wolf1728):

It is not B

OpenStudy (anonymous):

hmm... A?

OpenStudy (anonymous):

since it's the only other one with 9/2

OpenStudy (wolf1728):

A calculates correctly we KNOW perimeter = 21.294 So let's see what A equals = 27/2 + (9/2) * sqroot (3) = 13.5 + 7.794 = 21.294

OpenStudy (anonymous):

OHH!! so it is A! thank you so much... this really helped!!

OpenStudy (wolf1728):

Glad to help you out